# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 5.4 Performance of IDHGR Methods
The performance of IDHGR methods in the most used benchmarks (IsoGD (chalearn-lap), SHREC (SHREC), DHG (DHG-CVPRW), NVIDIA Hand Gesture or nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov), MSASL (MS-ASL), WLASL (WLASL), CSL500 (CSL500) and AUTSL (AUTSL)) is shown Table 3. For ISLR, Uni-Sign (UNISIGNli2025) surpasses relevant methods in terms of recognition accuracy, but requires powerful hardware to be executed fast, due to its large size (592.1M parameters, most of them contained on the language model). Thus, smaller but less accurate models like NLA-SLR (NLA_SLR), SignBERT+ (SignBERT_ICCV) and (Weichao-Zhao) are more reasonable solutions for real-world applications. Even though the multimodal SAM-SLR (SAM_SLR) and SAM-SLR-v2 (jiang2021sign-sam-slr-v2) are very accurate models for ISLR, they require a large number of modalities that are not fused effectively (using late fusion), which may be suboptimal. Pose-based methods for HCI and HRI including TD-GCN (TD-GCN), STST (STST) and DSTSA-GCN (DSTSA-CGN_cui2025dstsa) achieve high recognition rates with relatively low computational cost. Although, the computational complexity of pose estimation and the effect of its incorrect predictions in more challenging environments (e.g. fast movements) may reduce their overall performance. Appearance-based methods are not affected by off-the-shelf pose estimators but they comprise many more parameters (up to 251.4M (3D_CDC_NAS_TIP_2021)).
@@
STGCN-LSTM (STGCN-LSTM-10965662) 2025 - - - - - - - - - - - 95.2% -
(ISLR-STGCN-attention-LSTM-ozdemir2025cross) 2025 - - - - - - - - - - - - 86.43%
HAN-2S (HAN-liu2025han) 2025 0.94M - 95% 92.86% 92.71% 89.15% - - - - - - -
DSTSA-GCN (DSTSA-CGN_cui2025dstsa) 2025 8.08M - 97.74% 95.37% 95.04% 93.57% - - - - - - -
Uni-Sign (UNISIGNli2025) 2025 592.1M - - - - - - 78.16% 76.97% 63.52% 61.32% - -
TMS-Net (TMS-NET) 2024 - - - - - - - - - 56.4% - - -
MSNN (MSNN) 2024 - - - - - - - 50.59% - 47.26% - - -
SKIM (SKIM) 2024 - - - - - - - - - 57.73% 55.40% - -
MASA (MASA) 2024 - - - - - - - - - 55.77% 53.13% - -
DSTA-SLR (DSTA_SLR_hu2024dynamic) 2024 7.4M - - - - - - 65.74% 62.31% 53.68% 51.17% 98.1%5 -
(Weichao-Zhao) 2024 - - - - - - - 74.18% 72.18% 58.06% 55.66% 97.8% -
TD-GCN (TD-GCN) 2024 1.36M - 97.02% 95.36% 93.9% 91.4% - - - - - - -
STST (STST) 2024 - - 97.62% 95.83% 94.82% 93.18% - - - - - - -
(Tianyu_Liu) 2024 - - - - - - - - - - - 98.87% 92.53%
SwC GR-MMixer (SwC-GR-MMixer) 2024 - - - - - - - - - - - 98.54% -
HWGAT (FDMSE-hierarchicalwindowedgraphattention) 2024 - - - - - - - - - 48.49% - - 95.80%
(Yangke_patcog) 2024 0.21M - 96.9% 94.17% 94.21% 92.11% - - - - - - -
(omar_ikne) 2024 - - 94.1% 90.0% - - - - - - - - -
BEST (BEST) 2023 - - - - - - - 71.21% 68.24% 54.59% 52.12% 97.7% -
SignBERT+ (SignBERT+) 2023 - - - - - - - 73.71% 70.77% 55.59% 53.33% 97.8% -
NLA-SLR (NLA_SLR) 2023 - - - - - - - 73.80% 70.95% 61.26% 58.31% - -
3SCKI (3STREAM_DISTILLATION) 2023 - 65.87% - - - - - - - - - - -
(Yunan_Li) 2023 - 74.08% - - - - - - - - - - -
CTFB (CTFB_Hampiholi) 2023 27.2M-37.1M 69.02% - - - - 87.76% - - - - - -
L-TSSI-DenseNet (L-TSSI-DenseNet) 2023 7.2M - - - - - - - - - - - 93.13%
TS3C-Net (TS3C-Net) 2023 28.7M - - - - - 87.55% - - - - - -
(Musa_IEEE_ACCESS) 2023 - - 97.01% 92.78% 92.00% 88.78% - - - - - - -
SBI-DHGR (SBI-DHGR) 2023 28M - 97% 92.36% 94.64% 91.79% - - - - - - -
STr-GCN (STr-GCN) 2023 - - 93.39% 89.20% - - - - - - - - -
DRL-DDNet (KEYFRAMES_DRL_SKELETON) 2023 - - 96.5% 93.8% - - - - - - - - -
De+Recouple (Benjia_Zhou) 2022 - 66.79% - - - - 91.70% - - - - - -
MS-ISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) 2022 7.1M - 96.7% 94.9% 93.7% 91.2% - - - - - - -
SignBERT (SignBERT_ICCV) 2021 - - - - - - - 71.24% 67.96% 54.69% 52.08% 97.6% -
SlowFast+NAS+CDC (3D_CDC_NAS_TIP_2021) 2021 251.4M 66.23% - - - - - - - - - - -
RAAR3DNet (RAAR3DNet) 2021 - 66.62% - - - - 88.59% - - - - - -
SeST (SeST) 2021 - 60.27% - - - - - - - - - - -
SAM-SLR-v2 (jiang2021sign-sam-slr-v2) 2021 - - - - - - - - - 59.39% 56.63% 99.00% 98.53%
SAM-SLR (SAM_SLR) 2021 19.2M - - - - - - - - 58.73% 55.93% - 98.53%
(Hu_Zhou_Li_2021) 2021 - - - - - - - 69.39% 66.54% 51.39% 48.75% 98.3% -
@@ 6 Methods for Continuous Dynamic HGR
In contrast to IDHGR, where each sample contains at most one gesture, Continuous Hand Gesture Recognition (CHGR) deals with unsegmented gesture sequences. The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence. CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss (Fangyun_Wei_ICCV; CTC_graves2006connectionist).
While CHGR has been explored in domains like autonomous driving (FA-STGCN; CTCX_WANG2022123), most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss. CSLR aims to produce a sequence of glosses [^1] Glosses are atomic lexical units in sign language, analogous to words in spoken language. from video input. Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy. Some CSLR methods are further extended to support gloss-based Sign Language Translation (SLT) by adding a gloss-to-text module (MSKA_ELSEVIER; zuo2024onlinecontinuoussignlanguage; NEURIPS2022_6cd3ac24_HEATMAPS).
@@ 6.1 RGB-based
Most state-of-the-art RGB-based CHGR models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction (TemporalLiftPooling; VAC-CSLR; SMKD; CorrNet; C2ST; HezhenHu_TMM; SEN-CSLR; HezhenHu-JunfuPu; corrnetplus_hu2024corrnetsignlanguagerecognition). This standard architecture was introduced in (first_2dcnn_tcn_bilstm_TMM_2019). This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries (ADADOGLOU).
@@
Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives. For example, VAC‐CSLR (VAC-CSLR) augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig. 7). Similarly, SMKD (SMKD) adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR (C2SLR) enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
@@
Figure 7: An overview of the VAC-CSLR (VAC-CSLR). The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence. $F_{a}$ and $F_{b}$ denote auxiliary classifiers, while $L_{VE}$, $L_{VA}$ and $L_{CTC}$ the VE, VA and CTC loss, respectively. This training setting constitutes the baseline for most of the proposed methods.
@@
To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN. Temporal Lift Pooling (TLP) (TemporalLiftPooling) adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels. mLTSF‐Net (mLTSF-Net) couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales. HS‐I3D (HS-I3D) extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
@@
Other methods inject lightweight attention and correlation blocks into the spatial encoder. CorrNet (CorrNet) adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) further replaces its correlation operator with a more efficient variant and inserts an attention layer. SEN‐CSLR (SEN-CSLR) introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead. STNet (STNet-wang2025continuous) applied frame-wise loss to the 2DCNN backbone output to enhance feature extraction and inserted multiple Spatial Resonance Modules to facilitate inter-frame modeling.
@@
Many recent approaches replace 2DCNNs with transformer-based or hybrid CNN/attention modules. CLIP-SLA variants (CLIP_SLA_alyami2025clip_CSLR_2025) use a frozen CLIP visual encoder with LoRA or MLP adapters and an auxiliary classifier for early supervision. Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) combines a Swin transformer with an MS-TCN using TLP and dilated convolutions, followed by a BiLSTM for global context.
@@
TCNet (TCNet_Lu_Salah_Poppe_2024) augments a CNN backbone with trajectory and correlation modules, while SlowFastSign (SlowFastSign) modifies Slow-Fast-101 with bidirectional fusion and auxiliary temporal classifiers. To better model spatial correlations, SignGraph (SignGraph_CVPR) treats frame patches as nodes in a dynamic graph, alternating between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules; its multiscale variant, MultiSignGraph, further improves spatiotemporal context.
@@
CORE (CORE-CSLR-zhang2025core) proposed a different approach aiming to leverage the relations between regions within an input frame, deploying an off-the-shelf pose estimator to detect fine-grained and contextual cues; the corresponding frame region are fed into a 2DCNN and the extracted features are fused using a GNN to obtain the final embedding. These per-frame representations pass through a BiLSTM to perform global modeling. SD-Net (SD-Net-CSLR-gao2025structure) enhanced temporal modeling by using three parallel branches. Each branch consists of a two-stage temporal context module that captures both frame-wise and gloss-wise features. Contrastive loss is applied between two of the branches, based on the proposed Entity and Template scheme. Authors in (diffusion-CSLR-geng2026no) formulated CSLR as a cross-modal alignment task using a diffusion model to generate the glosses aiming to mitigate the weak supervision, achieving superior performance. DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) replaced the BiLSTM with a custom contextual model, consisting of memory and feed-forward blocks. Furthermore, the 2DCNN backbone was enhanced by adding a Spatial-Channel Attention (SCA) mechanism, supervised by ground-truth heatmaps, provided by an off-the-shelf pose estimator.
@@
Some models integrate linguistic embeddings with visual features to leverage gloss co-occurrence and language structure. SignBERT-HKSL (SignBERT-HKSL) combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and fine-tuning iteratively on masked outputs. C2ST (C2ST) replaces the CTC decoder with a Conditional Gloss Decoder (CGD) to model gloss dependencies and aligns visual and gloss embeddings via cross-modal match loss. GPGN (GPGN) similarly applies a cross-modal match loss between BERT-derived gloss features and PDC-TFE visual embeddings, followed by a BiGRU for sequence modeling.
@@
CVT-SLR (CVT-SLR) pretrains an asymmetric VAE on a pseudo gloss-to-gloss task, then connects it to a visual encoder with a video–gloss adapter and contrastive alignment loss. Fully transformer-based PiSLTRc (PiSLTRc) uses content-aware neighborhood gathering and position-informed attention for end-to-end CSLR, while (HezhenHu-JunfuPu) introduces prior-aware cross-modality augmentation, training a transformer to reconstruct masked gloss sequences and generate pseudo video–text pairs via guided gloss substitutions and insertions.
@@
Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations. Fangyun Wei et al. (Fangyun_Wei_ICCV) build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space. CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.

### v3

@@ 5.4 Performance of IDHGR Methods
The performance of IDHGR methods in the most used benchmarks (IsoGD (197), SHREC (185), DHG (39), NVIDIA Hand Gesture or nvGesture (138), MSASL (94), WLASL (110), CSL500 (87) and AUTSL (183)) is shown Table 3. For ISLR, Uni-Sign (113) surpasses relevant methods in terms of recognition accuracy, but requires powerful hardware to be executed fast, due to its large size (592.1M parameters, most of them contained on the language model). Thus, smaller but less accurate models like NLA-SLR (235), SignBERT+ (78) and (225) are more reasonable solutions for real-world applications. Even though the multimodal SAM-SLR (91) and SAM-SLR-v2 (90) are very accurate models for ISLR, they require a large number of modalities that are not fused effectively (using late fusion), which may be suboptimal. Pose-based methods for HCI and HRI including TD-GCN (120), STST (222) and DSTSA-GCN (32) achieve high recognition rates with relatively low computational cost. Although, the computational complexity of pose estimation and the effect of its incorrect predictions in more challenging environments (e.g. fast movements) may reduce their overall performance. Appearance-based methods are not affected by off-the-shelf pose estimators but they comprise many more parameters (up to 251.4M (215)).
@@
STGCN-LSTM (68) 2025 - - - - - - - - - - - 95.2% -
(149) 2025 - - - - - - - - - - - - 86.43%
HAN-2S (119) 2025 0.94M - 95% 92.86% 92.71% 89.15% - - - - - - -
DSTSA-GCN (32) 2025 8.08M - 97.74% 95.37% 95.04% 93.57% - - - - - - -
Uni-Sign (113) 2025 592.1M - - - - - - 78.16% 76.97% 63.52% 61.32% - -
TMS-Net (40) 2024 - - - - - - - - - 56.4% - - -
MSNN (130) 2024 - - - - - - - 50.59% - 47.26% - - -
SKIM (115) 2024 - - - - - - - - - 57.73% 55.40% - -
MASA (223) 2024 - - - - - - - - - 55.77% 53.13% - -
DSTA-SLR (86) 2024 7.4M - - - - - - 65.74% 62.31% 53.68% 51.17% 98.1%5 -
(225) 2024 - - - - - - - 74.18% 72.18% 58.06% 55.66% 97.8% -
TD-GCN (120) 2024 1.36M - 97.02% 95.36% 93.9% 91.4% - - - - - - -
STST (222) 2024 - - 97.62% 95.83% 94.82% 93.18% - - - - - - -
(123) 2024 - - - - - - - - - - - 98.87% 92.53%
SwC GR-MMixer (122) 2024 - - - - - - - - - - - 98.54% -
HWGAT (152) 2024 - - - - - - - - - 48.49% - - 95.80%
(111) 2024 0.21M - 96.9% 94.17% 94.21% 92.11% - - - - - - -
(88) 2024 - - 94.1% 90.0% - - - - - - - - -
BEST (224) 2023 - - - - - - - 71.21% 68.24% 54.59% 52.12% 97.7% -
SignBERT+ (77) 2023 - - - - - - - 73.71% 70.77% 55.59% 53.33% 97.8% -
NLA-SLR (235) 2023 - - - - - - - 73.80% 70.95% 61.26% 58.31% - -
3SCKI (198) 2023 - 65.87% - - - - - - - - - - -
(112) 2023 - 74.08% - - - - - - - - - - -
CTFB (66) 2023 27.2M-37.1M 69.02% - - - - 87.76% - - - - - -
L-TSSI-DenseNet (108) 2023 7.2M - - - - - - - - - - - 93.13%
TS3C-Net (70) 2023 28.7M - - - - - 87.55% - - - - - -
(135) 2023 - - 97.01% 92.78% 92.00% 88.78% - - - - - - -
SBI-DHGR (143) 2023 28M - 97% 92.36% 94.64% 91.79% - - - - - - -
STr-GCN (184) 2023 - - 93.39% 89.20% - - - - - - - - -
DRL-DDNet (51) 2023 - - 96.5% 93.8% - - - - - - - - -
De+Recouple (228) 2022 - 66.79% - - - - 91.70% - - - - - -
MS-ISTGCN (186) 2022 7.1M - 96.7% 94.9% 93.7% 91.2% - - - - - - -
SignBERT (78) 2021 - - - - - - - 71.24% 67.96% 54.69% 52.08% 97.6% -
SlowFast+NAS+CDC (215) 2021 251.4M 66.23% - - - - - - - - - - -
RAAR3DNet (227) 2021 - 66.62% - - - - 88.59% - - - - - -
SeST (191) 2021 - 60.27% - - - - - - - - - - -
SAM-SLR-v2 (90) 2021 - - - - - - - - - 59.39% 56.63% 99.00% 98.53%
SAM-SLR (91) 2021 19.2M - - - - - - - - 58.73% 55.93% - 98.53%
(79) 2021 - - - - - - - 69.39% 66.54% 51.39% 48.75% 98.3% -
@@ 6 Methods for Continuous Dynamic HGR
In contrast to IDHGR, where each sample contains at most one gesture, Continuous Hand Gesture Recognition (CHGR) deals with unsegmented gesture sequences. The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence. CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss (207; 61).
While CHGR has been explored in domains like autonomous driving (50; 204), most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss. CSLR aims to produce a sequence of glosses [^1] Glosses are atomic lexical units in sign language, analogous to words in spoken language. from video input. Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy. Some CSLR methods are further extended to support gloss-based Sign Language Translation (SLT) by adding a gloss-to-text module (62; 236; 29).
@@ 6.1 RGB-based
Most state-of-the-art RGB-based CHGR models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction (83; 136; 69; 84; 217; 76; 85; 75; 82). This standard architecture was introduced in (33). This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries (2).
@@
Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives. For example, VAC‐CSLR (136) augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig. 7). Similarly, SMKD (69) adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR (234) enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
@@
Figure 7: An overview of the VAC-CSLR (136). The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence. $F_{a}$ and $F_{b}$ denote auxiliary classifiers, while $L_{VE}$, $L_{VA}$ and $L_{CTC}$ the VE, VA and CTC loss, respectively. This training setting constitutes the baseline for most of the proposed methods.
@@
To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN. Temporal Lift Pooling (TLP) (83) adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels. mLTSF‐Net (210) couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales. HS‐I3D (209) extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
@@
Other methods inject lightweight attention and correlation blocks into the spatial encoder. CorrNet (84) adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ (82) further replaces its correlation operator with a more efficient variant and inserts an attention layer. SEN‐CSLR (85) introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead. STNet (205) applied frame-wise loss to the 2DCNN backbone output to enhance feature extraction and inserted multiple Spatial Resonance Modules to facilitate inter-frame modeling.
@@
Many recent approaches replace 2DCNNs with transformer-based or hybrid CNN/attention modules. CLIP-SLA variants (14) use a frozen CLIP visual encoder with LoRA or MLP adapters and an auxiliary classifier for early supervision. Swin-MSTP (15) combines a Swin transformer with an MS-TCN using TLP and dilated convolutions, followed by a BiLSTM for global context.
@@
TCNet (124) augments a CNN backbone with trajectory and correlation modules, while SlowFastSign (4) modifies Slow-Fast-101 with bidirectional fusion and auxiliary temporal classifiers. To better model spatial correlations, SignGraph (52) treats frame patches as nodes in a dynamic graph, alternating between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules; its multiscale variant, MultiSignGraph, further improves spatiotemporal context.
@@
CORE (219) proposed a different approach aiming to leverage the relations between regions within an input frame, deploying an off-the-shelf pose estimator to detect fine-grained and contextual cues; the corresponding frame region are fed into a 2DCNN and the extracted features are fused using a GNN to obtain the final embedding. These per-frame representations pass through a BiLSTM to perform global modeling. SD-Net (53) enhanced temporal modeling by using three parallel branches. Each branch consists of a two-stage temporal context module that captures both frame-wise and gloss-wise features. Contrastive loss is applied between two of the branches, based on the proposed Entity and Template scheme. Authors in (59) formulated CSLR as a cross-modal alignment task using a diffusion model to generate the glosses aiming to mitigate the weak supervision, achieving superior performance. DSTEN-CSLR (214) replaced the BiLSTM with a custom contextual model, consisting of memory and feed-forward blocks. Furthermore, the 2DCNN backbone was enhanced by adding a Spatial-Channel Attention (SCA) mechanism, supervised by ground-truth heatmaps, provided by an off-the-shelf pose estimator.
@@
Some models integrate linguistic embeddings with visual features to leverage gloss co-occurrence and language structure. SignBERT-HKSL (233) combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and fine-tuning iteratively on masked outputs. C2ST (217) replaces the CTC decoder with a Conditional Gloss Decoder (CGD) to model gloss dependencies and aligns visual and gloss embeddings via cross-modal match loss. GPGN (64) similarly applies a cross-modal match loss between BERT-derived gloss features and PDC-TFE visual embeddings, followed by a BiGRU for sequence modeling.
@@
CVT-SLR (226) pretrains an asymmetric VAE on a pseudo gloss-to-gloss task, then connects it to a visual encoder with a video–gloss adapter and contrastive alignment loss. Fully transformer-based PiSLTRc (211) uses content-aware neighborhood gathering and position-informed attention for end-to-end CSLR, while (75) introduces prior-aware cross-modality augmentation, training a transformer to reconstruct masked gloss sequences and generate pseudo video–text pairs via guided gloss substitutions and insertions.
@@
Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations. Fangyun Wei et al. (207) build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space. CSGC (161) maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.